Age is something I am proud of said grandfather. "I would be as old now as six times my age six years from now less six times my age six years ago if I were six years older than I am." How old am I?
Answer by ankor@dixie-net.com(22740)

Age is something I am proud of said grandfather.
"I would be as old now as six times my age six years from now less six times my age six years ago if I were six years older than I am."
:
let a = his age now
a + 6 = 6(a+6) - 6(a-6)
a + 6 = 6a + 36 - 6a + 36
a = 36 + 36 - 6
a = 72 - 6
a = 66 yrs old
:
:
Check this
66 + 6 = 6(66+6) - 6(66-6)
72 = 432 - 360